OpenStudy (delina):

a current of 4 A flows from a battery when a light bulb is connected across its terminals. The P.D across the terminal is 12V. a) what quantity of charge leaves the battery every second. b) how much potential energy does each coulomb leaving the battery possess. c) how much charge must be given out by the battery if it is to supply 60 joules of energy? d) How long does the battery take to supply 60 joules of energy? I will fan and vote for the people that answer

OpenStudy (joseph127):

a). Q=IT Q=12*1. Q=12columbs b). P.E = W W=QV. W=4*12. W=48J c). W=QV. 60=Q*12 Q=60/12 Q=5columbs d). Q=IT 60=4*T T=60/4 T= 15sec Q=quantity of charge I=current T=time W=work